# From Restless to Contextual: A Thresholding Bandit Approach to Improve Finite-horizon Performance

### v2

@@ Appendix C Technical Details In Section 3
$\Rightarrow$: When Assumption 3.1 holds, suppose there exists $s\in\mathcal{S}$ such that $q_{s1}^{1}<q_{s1}^{0}$. Without loss of generality (WLOG), assume $s=1$. We have for any increasing function $f$:
$\mathbb{E}_{s^{\prime}\sim q_{1s^{\prime}}^{1}}\left[f(s^{\prime})\right]-\mathbb{E}_{s^{\prime}\sim q_{1s^{\prime}}^{0}}\left[f(s^{\prime})\right]$ $=f(1)q^{1}_{11}+f(0)q^{1}_{10}-f(1)q^{0}_{11}-f(0)q^{0}_{10}$
$=f(1)\left(q^{1}_{11}-q^{0}_{11}\right)+f(0)\left(q^{1}_{10}-q^{0}_{10}\right)$
$=\left(f(1)-f(0)\right)\left(q^{1}_{11}-q^{0}_{11}\right)$
$<0,$
where the last inequality holds because $f$ is increasing and the assumption that $q^{1}_{11}<q^{0}_{11}$. This will contradict Assumption 3.1. Therefore, we must have for all $s\in\mathcal{S}$, $q_{s1}^{1}\geq q_{s1}^{0}$.
:absent:$\Leftarrow$:⇐ : When for all $s\in\mathcal{S}$, $q_{s1}^{1}\geq q_{s1}^{0}$, we will prove Eq. (7) separately for $s=0$ and $s=1$. When $s=1$, we have
$\mathbb{E}_{s^{\prime}\sim q_{1s^{\prime}}^{1}}\left[f(s^{\prime})\right]-\mathbb{E}_{s^{\prime}\sim q_{1s^{\prime}}^{0}}\left[f(s^{\prime})\right]$ $=f(1)q^{1}_{11}+f(0)q^{1}_{10}-f(1)q^{0}_{11}-f(0)q^{0}_{10}$
$=f(1)\left(q^{1}_{11}-q^{0}_{11}\right)+f(0)\left(q^{1}_{10}-q^{0}_{10}\right)$
$=\left(f(1)-f(0)\right)\left(q^{1}_{11}-q^{0}_{11}\right)$
$\geq 0,$
where the last inequality holds because $f$ is increasing and the assumption that $q^{1}_{11}\geq q^{0}_{11}$. Similarly, we can show that $\mathbb{E}_{s^{\prime}\sim q_{0s^{\prime}}^{1}}\left[f(s^{\prime})\right]\geq\mathbb{E}_{s^{\prime}\sim q_{0s^{\prime}}^{0}}\left[f(s^{\prime})\right]$. This completes the proof. ∎
C.2 Proof of Lemma 3.3
See 3.3
Proof.
We will use Bellman optimality equation to prove the theorem. Specifically, we only need to prove that for any $h$, the maximizer $a$ of the following problem,
$V_{h}^{*}(s)=R(s)+\max_{a}\sum_{s^{\prime}}q_{ss^{\prime}}^{a}V_{h+1}^{*}(s),$
aligns with $\pi_{h+1}^{\text{greedy}}(s)$. With Assumption 3.1, this is equivalent to showing $V_{h+1}^{*}(s)$ is an increasing function in $s$. To do this, we will discuss two cases: 1) $q_{11}^{1}\geq q_{01}^{1}$ and 2) $q_{11}^{1}\leq q_{01}^{1}$.
Case 1:
$q^{1}_{11}\geq q^{1}_{01}$. We will use backward induction to prove the following statements:
$\forall t\leq T,~{}$ $V^{*}_{t}(1)\geq V_{t}^{*}(0);$ (C.16)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.17)
When $t=T$, the statements hold trivially. Suppose the statements holds for all $t\geq h+1$, when $t=h$, we will first prove that the greedy policy is optimal assuming $V^{*}_{h+1}(1)\geq V_{h+1}^{*}(0)$. By Bellman optimality equation, we have
$V_{h}^{*}(s)=R(s)+\max_{a}\sum_{s^{\prime}}q^{a}_{ss^{\prime}}V_{h+1}^{*}\left(s^{\prime}\right).$
Since $V_{h+1}^{*}$ is increasing function, by Assumption 3.1, we have
$\sum_{s^{\prime}}q^{1}_{ss^{\prime}}V_{h+1}^{*}\left(s^{\prime}\right)\geq\sum_{s^{\prime}}q^{0}_{ss^{\prime}}V_{h+1}^{*}\left(s^{\prime}\right).$
This implies that action 1 is the maximizer, and the optimal policy will choose to give intervention at $t$, indicating Eq. (C.17) holds. Next, assuming that $\pi_{h+1}^{\text{greedy}}=\pi_{h+1}^{*}$, we prove for $V^{*}_{h}(1)\geq V_{h}^{*}(0)$:
$V_{h}^{*}(1)-V_{h}^{*}(0)$ $=R(1)-R(0)+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{1,s^{\prime}}V_{h+1}^{*}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{0,s^{\prime}}V_{h+1}^{*}(s^{\prime})$
$\geq\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}(1)-V_{h+1}^{*}(0)\right)$
$\geq 0.$
Case 2:
$q^{1}_{11}\leq q^{1}_{01}$. Similarly, we use backward induction to prove the following statements:
$\forall t\leq T,~{}$ $0\leq V^{*}_{t}(1)-V_{t}^{*}(0)\leq R(1)-R(0);$ (C.18)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.19)
This trivially holds for $t=T$. Suppose this holds for $t\geq h+1$, when $t=h$, we can prove Eq. (C.19) for $t=h$ using exactly the same logic. Assuming that $\pi_{h+1}^{\text{greedy}}=\pi_{h+1}^{*}$ and that Eq. (C.18) hold for $t=h+1$, we have
$V_{h}^{*}(1)-V_{h}^{*}(0)$ $=R(1)-R(0)+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{1,s^{\prime}}V_{h+1}^{*}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{0,s^{\prime}}V_{h+1}^{*}(s^{\prime})$
$=R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}(1)-V_{h+1}^{*}(0)\right)$
$\leq R(1)-R(0).$
This will give us the RHS of Eq. (C.18) for $t=h$. Next, we will show the LHS of Eq. (C.18) for $t=h$ under the same assumption that $\pi_{h+1}^{\text{greedy}}=\pi_{h+1}^{*}$ and that Eq. (C.18) hold for $t=h+1$.
$V_{h}^{*}(1)-V_{h}^{*}(0)$ $=R(1)-R(0)+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{1,s^{\prime}}V_{h+1}^{*}(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{0,s^{\prime}}V_{h+1}^{*}(s^{\prime})$
$=R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}(1)-V_{h+1}^{*}(0)\right)$
$\geq R(1)-R(0)+\left(q^{1}_{11}-q^{1}_{01}\right)\left(R(1)-R(0)\right)$
$=\left(q^{1}_{11}+q_{00}^{1}\right)\left(R(1)-R(0)\right)$
$\geq 0.$
Together, they will give us Eq. (C.18) for $t=h$. This completes the proof. ∎
C.3 Proof of Lemma 3.4
See 3.4
Proof.
WLOG, assume $M=2,B=1$. Again, using Bellman optimal equation, we only need to prove that for any $t\leq T$, the following two equations hold for any $\left(s_{t}^{1},s_{t}^{2}\right)\in\mathcal{S}^{2}$:
$\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{1}_{s_{t}^{1},s_{t+1}^{1}}(1)q^{0}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right)$
$\geq\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{0}_{s_{t}^{1}s_{t+1}^{1}}(1)q^{0}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right)$ (C.20)
$\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{0}_{s_{t}^{1}s_{t+1}^{1}}(1)q^{1}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right)$
$\geq\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{0}_{s_{t}^{1}s_{t+1}^{1}}(1)q^{0}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right).$ (C.21)
It is easy to see that these two equations are symmetric, thus, we will only provide the proof for the first inequality. Same as the proof of Lemma 3.3, we will prove the result under two cases: 1) $q^{1}_{11}(1)\geq q^{1}_{01}(1)$ and 2) $q^{1}_{11}(1)\leq q^{1}_{01}(1)$.
Case 1:
$q^{1}_{11}(1)\geq q^{1}_{01}(1)$. We will use backward induction to prove the following statements:
$\forall t\leq T,~{}$ $f_{s,t}(1)\geq f_{s,t}(0),\forall s\in\mathcal{S};$ (C.22)
$\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{1}_{s_{t}^{1},s_{t+1}^{1}}(1)q^{0}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right)$
$\geq\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{0}_{s_{t}^{1}s_{t+1}^{1}}(1)q^{0}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right),\forall s_{t}^{i}\in\mathcal{S},$ (C.23)
where $f_{s,t}\left(s^{\prime}\right):=\sum_{s_{t+1}^{2}\in\mathcal{S}}q^{0}_{ss_{t+1}^{2}}\left(2\right)V_{t+1}^{*}\left(\left(s^{\prime},s_{t+1}^{2}\right)\right)$. They hold trivially when $t=T$. Suppose the statements hold for $t\geq h+1$, when $t=h$, we will first prove Eq. (C.22) for $t=h$ assuming that Eq.(C.22) holds for $t=h+1$.
$f_{s,h}(1)-f_{s,h}(0)$ $=\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{t+1}^{2}}\left(2\right)V_{h+1}^{*}\left(\left(1,s_{h+1}^{2}\right)\right)-\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{t+1}^{2}}\left(2\right)V_{h+1}^{*}\left(\left(0,s_{h+1}^{2}\right)\right)$
$=\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(R^{1}(1)+R^{2}(s_{h+1}^{2})+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{1s^{\prime}}(1)f_{s_{h+1}^{2},h+1}(s^{\prime})\right)$
$-\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(R^{1}(0)+R^{2}(s_{h+1}^{2})+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{0s^{\prime}}(1)f_{s_{h+1}^{2},h+1}(s^{\prime})\right)$
$=R^{1}(1)-R^{1}(0)+\left(q^{1}_{11}(1)-q^{1}_{01}(1)\right)\left(\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(f_{s_{h+1}^{2},h+1}(1)-f_{s_{h+1}^{2},h+1}(0)\right)\right)$
$\geq 0.$
Next, we want to show Eq. (C.23) for $t=h$ assuming that Eq. (C.22) holds for $t=h$.
$\sum_{\left(s_{h+1}^{1},s_{h+1}^{2}\right)\in\mathcal{S}^{2}}q^{1}_{s_{h}^{1}s_{h+1}^{1}}(1)q^{0}_{s_{h}^{2}s_{h+1}^{2}}(2)V_{h+1}^{*}\left(\left(s_{h+1}^{1},s_{h+1}^{2}\right)\right)$
$-\sum_{\left(s_{h+1}^{1},s_{h+1}^{2}\right)\in\mathcal{S}^{2}}q^{0}_{s_{h}^{1}s_{h+1}^{1}}(1)q^{0}_{s_{h}^{2}s_{h+1}^{2}}(2)V_{h+1}^{*}\left(\left(s_{h+1}^{1},s_{h+1}^{2}\right)\right)$
$=\sum_{s_{h+1}^{1}}q^{1}_{s_{h}^{1}s_{h+1}^{1}}(1)f_{s_{h}^{2},h}\left(s_{h+1}^{1}\right)-\sum_{s_{h+1}^{1}}q^{0}_{s_{h}^{1}s_{h+1}^{1}}(1)f_{s_{h}^{2},h}\left(s_{h+1}^{1}\right)$
$\geq 0,$
where the last inequality holds because of Eq. (C.22) and Assumption 3.1.
Case 2:
$q^{1}_{11}(1)\leq q^{1}_{01}(1)$. We will use the backward induction to prove the following statements:
$\forall t\leq T,~{}$ $0\leq f_{s,t}(1)-f_{s,t}(0)\leq R^{1}(1)-R^{1}(0),\forall s\in\mathcal{S};$ (C.24)
$\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{1}_{s_{t}^{1},s_{t+1}^{1}}(1)q^{0}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right)$
$\geq\sum_{\left(s_{t+1}^{1},s_{t+1}^{2}\right)\in\mathcal{S}^{2}}q^{0}_{s_{t}^{1}s_{t+1}^{1}}(1)q^{0}_{s_{t}^{2}s_{t+1}^{2}}(2)V_{t+1}^{*}\left(\left(s_{t+1}^{1},s_{t+1}^{2}\right)\right),\forall s_{t}^{i}\in\mathcal{S}.$ (C.25)
They trivially hold for $t=T$. Suppose the statements hold for $t\geq h+1$, when $t=h$, we will first prove the RHS of Eq. (C.24) for $t=h$ assuming that Eq. (C.24) holds for $t=h+1$.
$\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)V_{h+1}^{*}\left(\left(1,s_{h+1}^{2}\right)\right)-\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)V_{h+1}^{*}\left(\left(0,s_{h+1}^{2}\right)\right)$
$=\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(R^{1}(1)+R^{2}(s_{h+1}^{2})+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{1s^{\prime}}(1)f_{s_{h+1}^{2},h+1}(s^{\prime})\right)$
$-\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(R^{1}(0)+R^{2}(s_{h+1}^{2})+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{0s^{\prime}}(1)f_{s_{h+1}^{2},h+1}(s^{\prime})\right)$
$=R^{1}(1)-R^{1}(0)+\left(q^{1}_{11}(1)-q^{1}_{01}(1)\right)\left(\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(f_{s_{h+1}^{2},h+1}(1)-f_{s_{h+1}^{2},h+1}(0)\right)\right)$
$\leq R^{1}(1)-R^{1}(0).$
Then we want to prove the LHS of Eq. (C.24) under the same assumption.
$\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)V_{h+1}^{*}\left(\left(1,s_{h+1}^{2}\right)\right)-\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)V_{h+1}^{*}\left(\left(0,s_{h+1}^{2}\right)\right)$
$=\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(R^{1}(1)+R^{2}(s_{h+1}^{2})+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{1s^{\prime}}(1)f_{s_{h+1}^{2},h+1}(s^{\prime})\right)$
$-\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(R^{1}(0)+R^{2}(s_{h+1}^{2})+\sum_{s^{\prime}\in\mathcal{S}}q^{1}_{0s^{\prime}}(1)f_{s_{h+1}^{2},h+1}(s^{\prime})\right)$
$=R^{1}(1)-R^{1}(0)+\left(q^{1}_{11}(1)-q^{1}_{01}(1)\right)\left(\sum_{s_{h+1}^{2}\in\mathcal{S}}q^{0}_{ss_{h+1}^{2}}(2)\left(f_{s_{h+1}^{2},h+1}(1)-f_{s_{h+1}^{2},h+1}(0)\right)\right)$
$\geq R^{1}(1)-R^{1}(0)+\left(q^{1}_{11}(1)-q^{1}_{01}(1)\right)\left(R^{1}(1)-R^{1}(0)\right)$
$\geq 0.$
This shows the LHS of Eq. (C.24). Using the same logic used in Case 1 will give us Eq. (C.25) for $t=h$. ∎
C.4 Proof of Theorem 3.6
See 3.6
Proof.
By Lemma 3.4, we only need to care which $B$ agents do we need to give action to. WLOG, assume that $I(1)\leq I(0)$. For simplicity, we first look at the simple case where $M=2,B=1$. The generalization to multi-agent case is straightforward and will be deferred to the last part of the proof. First of all, we will derive some useful equations that are useful in the proof.
@@
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right).$ (C.26)
@@
$+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right).$ (C.27)
@@
$+q^{1}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right).$ (C.28)
@@
$=q^{1}_{11}R(1)+q^{1}_{10}R(0)-q^{0}_{11}R(1)-q^{0}_{10}R(0)$ (C.29)
$=\left(q^{1}_{11}-q^{0}_{11}\right)\left(R(1)-R(0)\right).$ (C.30)
@@
$I(0)=\left(q^{1}_{01}-q^{0}_{01}\right)\left(R(1)-R(0)\right).$ (C.31)
In the following, we will consider three cases.
@@
$q^{0}_{11}\geq q^{0}_{01},q^{1}_{11}\leq q^{1}_{01}$. We will use backward induction to prove the following statements:
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$ (C.32)
$0\leq V_{t}^{*}\left((1,1)\right)-V_{t}^{*}((1,0))\leq R(1)-R(0)$ (C.33)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.34)
The statements hold trivially when $t=T$. Suppose the statements hold when $t\geq h+1$, when $t=h$, we will first prove Eq. (C.34) for $t=h$. Recall that the policy satisfying the following Bellman optimality equation is the optimal policy:
$V_{h}^{*}\left((s_{h}^{1},s_{h}^{2})\right)=R(s_{h}^{1})+R(s_{h}^{2})+\max_{a_{h+1}^{1},a_{h+1}^{2}}\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{a_{h+1}^{1}}_{s_{h}^{1}s_{h+1}^{1}}q^{a_{h+1}^{2}}_{s_{h}^{2}s_{h+1}^{2}}V_{h+1}^{*}\left(s_{h+1}^{1},s_{h+1}^{2}\right).$ (C.35)
By Lemma 3.4, the maximizer is attained when one of the $a_{h+1}^{i}=1$. Therefore, it’s easy to see that when $s_{h}^{1}=s_{h}^{2}$, randomly giving action will satisfy this equation. We will focus on the case where $s_{h}^{1}\neq s_{h}^{2}$. In that case, to show Eq. (C.34), we only need to show that
$\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{0}_{1s_{h+1}^{1}}q^{1}_{0s_{h+1}^{2}}V_{h+1}^{*}\left(s_{h+1}^{1},s_{h+1}^{2}\right)\geq\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{1}_{1s_{h+1}^{1}}q^{0}_{0s_{h+1}^{2}}V_{h+1}^{*}\left(s_{h+1}^{1},s_{h+1}^{2}\right).$ (C.36)
By Eq. (C.26), this is equivalent to showing
@@
where the first inequality holds because of the induction hypothesis and $q^{0}_{11}\geq q^{0}_{01},q^{1}_{01}\geq q^{1}_{11}$, the first equality holds because of Eq. (C.31) and Eq. (C.30) and the last inequality holds because $q^{1}_{10}\geq q^{1}_{00}$, $I(0)\geq I(1)$ and the assumption that Eq. (C.32) holds for $t=h+1$.
Next, we first show the LHS of Eq. (C.32) for $t=h$ under the assumption that Eq. (C.34) holds for $t=h$ and Eq. (C.32) holds for $t=h+1$.
@@
where the first equality holds because of Eq. (C.28) and the last inequality holds since $q^{0}_{11}\geq q^{0}_{01}$. Under the same condition, for the RHS of Eq. (C.32), we have
@@
Then, we only need to show Eq. (C.33) for $t=h$. Assuming that Eq. (C.34) holds for $t=h$ and Eq. (C.33) holds for $t=h+1$, for the RHS, we have
@@
where the last inequality holds since $q^{1}_{11}\geq q^{0}_{11}$ by Lemma 3.2.
@@
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))\leq R(1)-R(0)$ (C.37)
$0\leq V_{t}^{*}\left((1,1)\right)-V_{t}^{*}((1,0))\leq R(1)-R(0)$ (C.38)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.39)
When $t=T$, the statements hold trivially. Suppose the statements hold for $t\geq h+1$, when $t=h$, we will first prove Eq. (C.39) under the assumption that Eq. (C.37) and Eq. (C.38) holds for $t=h+1$. Similarly we only need to show Eq. (C.36).
@@
where the first inequality holds because the assumption we make and the last inequality holds because $q^{1}_{10}\geq q^{1}_{00}$, $I(0)\geq I(1)$ and the assumption that Eq. (C.37) holds.
Next we want to prove the RHS of Eq. (C.38) under the assumption that Eq. (C.39) holds for $t=h$ and that Eq. (C.37) and Eq. (C.38) hold for $t=h+1$
@@
Under the same condition, for the LHS of Eq. (C.38), we have
@@
Similarly, we can show Eq. (C.37) for $t=h$ under the same condition.
@@
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))\leq{R(1)-R(0)}$ (C.40)
$0\leq V_{t}^{*}\left((1,1)\right)-V_{t}^{*}((1,0))\leq R(1)-R(0)$ (C.41)
$V_{t}^{*}\left((1,1)\right)-V_{t}^{*}((1,0))\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (C.42)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.43)
When $t=T$, the statements hold trivially. Suppose this holds for $t\geq h+1$, when $t=h$, we will first show Eq. (C.43) for $t=h$ under the assumption that Eq. (C.40), Eq. (C.41) and Eq. (C.42) hold for $t=h+1$. Similarly, this is equivalent to show Eq. (C.36).
@@
where the first equality holds because of Eq. (C.31) and Eq. (C.30) and the first inequality holds since $q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\leq 0$ and the assumption that Eq. (C.42) holds for $t=h+1$.

### v4

@@ Appendix E Technical Details In Section 3
WLOG, assume that $I(1)\leq I(0)$. For simplicity, we first look at the simple case where $M=2,B=1$. The generalization to multi-agent case is straightforward and will be deferred to the last part of the proof. First of all, we will derive some useful equations that are useful in the proof.
@@
$+\left(q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right).$ (E.12)
@@
$+q^{0}_{11}\left(q^{1}_{11}-q^{1}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right).$ (E.13)
@@
$+q^{1}_{01}\left(q^{0}_{11}-q^{0}_{01}\right)\left(V_{h+1}^{*}\left((1,1)\right)-2V_{h+1}^{*}\left((1,0)\right)+V_{h+1}^{*}\left((0,0)\right)\right).$ (E.14)
@@
$=q^{1}_{11}R(1)+q^{1}_{10}R(0)-q^{0}_{11}R(1)-q^{0}_{10}R(0)$ (E.15)
$=\left(q^{1}_{11}-q^{0}_{11}\right)\left(R(1)-R(0)\right).$ (E.16)
@@
$I(0)=\left(q^{1}_{01}-q^{0}_{01}\right)\left(R(1)-R(0)\right).$ (E.17)
In the following, we will consider four cases.
@@
$q^{0}_{11}\geq q^{0}_{01},q^{1}_{11}\leq q^{1}_{01},q_{11}^{1}\geq q_{11}^{0}$. We will use backward induction to prove the following statements:
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))\leq\frac{R(1)-R(0)}{q^{0}_{10}}$ (E.18)
$0\leq V_{t}^{*}\left((1,1)\right)-V_{t}^{*}((1,0))\leq R(1)-R(0)$ (E.19)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (E.20)
The statements hold trivially when $t=T$. Suppose the statements hold when $t\geq h+1$, when $t=h$, we will first prove Eq. (E.20) for $t=h$. Recall that the policy satisfying the following Bellman optimality equation is the optimal policy:
$V_{h}^{*}\left((s_{h}^{1},s_{h}^{2})\right)=R(s_{h}^{1})+R(s_{h}^{2})+\max_{a_{h+1}^{1},a_{h+1}^{2}}\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{a_{h+1}^{1}}_{s_{h}^{1}s_{h+1}^{1}}q^{a_{h+1}^{2}}_{s_{h}^{2}s_{h+1}^{2}}V_{h+1}^{*}\left(s_{h+1}^{1},s_{h+1}^{2}\right).$ (E.21)
By the problem formulation, one of the $a_{h+1}^{i}=1$. Therefore, it’s easy to see that when $s_{h}^{1}=s_{h}^{2}$, randomly giving action will satisfy this equation. We will focus on the case where $s_{h}^{1}\neq s_{h}^{2}$. In that case, to show Eq. (E.20), we only need to show that
$\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{0}_{1s_{h+1}^{1}}q^{1}_{0s_{h+1}^{2}}V_{h+1}^{*}\left(s_{h+1}^{1},s_{h+1}^{2}\right)\geq\sum_{s_{h+1}^{1},s_{h+1}^{2}}q^{1}_{1s_{h+1}^{1}}q^{0}_{0s_{h+1}^{2}}V_{h+1}^{*}\left(s_{h+1}^{1},s_{h+1}^{2}\right).$ (E.22)
By Eq. (E.12), this is equivalent to showing
@@
where the first inequality holds because of the induction hypothesis and $q^{0}_{11}\geq q^{0}_{01},q^{1}_{01}\geq q^{1}_{11}$, the first equality holds because of Eq. (E.17) and Eq. (E.16) and the last inequality holds because $q^{1}_{10}\geq q^{1}_{00}$, $I(0)\geq I(1)$ and the assumption that Eq. (E.18) holds for $t=h+1$.
Next, we first show the LHS of Eq. (E.18) for $t=h$ under the assumption that Eq. (E.20) holds for $t=h$ and Eq. (E.18) holds for $t=h+1$.
@@
where the first equality holds because of Eq. (E.14) and the last inequality holds since $q^{0}_{11}\geq q^{0}_{01}$. Under the same condition, for the RHS of Eq. (E.18), we have
@@
Then, we only need to show Eq. (E.19) for $t=h$. Assuming that Eq. (E.20) holds for $t=h$ and Eq. (E.19) holds for $t=h+1$, for the RHS, we have
@@
where the last inequality holds since $q^{1}_{11}\geq q^{0}_{11}$.
@@
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))\leq R(1)-R(0)$ (E.23)
$0\leq V_{t}^{*}\left((1,1)\right)-V_{t}^{*}((1,0))\leq R(1)-R(0)$ (E.24)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (E.25)
When $t=T$, the statements hold trivially. Suppose the statements hold for $t\geq h+1$, when $t=h$, we will first prove Eq. (E.25) under the assumption that Eq. (E.23) and Eq. (E.24) holds for $t=h+1$. Similarly we only need to show Eq. (E.22).
@@
where the first inequality holds because the assumption we make and the last inequality holds because $q^{1}_{10}\geq q^{1}_{00}$, $I(0)\geq I(1)$ and the assumption that Eq. (E.23) holds.
Next we want to prove the RHS of Eq. (E.24) under the assumption that Eq. (E.25) holds for $t=h$ and that Eq. (E.23) and Eq. (E.24) hold for $t=h+1$ .
@@
Under the same condition, for the LHS of Eq. (E.24), we have
@@
Similarly, we can show Eq. (E.23) for $t=h$ under the same condition.
@@
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))\leq{R(1)-R(0)}$ (E.26)
$0\leq V_{t}^{*}\left((1,1)\right)-V_{t}^{*}((1,0))\leq R(1)-R(0)$ (E.27)
$V_{t}^{*}\left((1,1)\right)-V_{t}^{*}((1,0))\leq V_{t}^{*}\left((1,0)\right)-V_{t}^{*}((0,0))$ (E.28)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (E.29)
When $t=T$, the statements hold trivially. Suppose this holds for $t\geq h+1$, when $t=h$, we will first show Eq. (E.29) for $t=h$ under the assumption that Eq. (E.26), Eq. (E.27) and Eq. (E.28) hold for $t=h+1$. Similarly, this is equivalent to show Eq. (E.22).
@@
where the first equality holds because of Eq. (E.17) and Eq. (E.16) and the first inequality holds since $q^{0}_{11}q^{1}_{01}-q^{0}_{01}q^{1}_{11}\leq 0$ and the assumption that Eq. (E.28) holds for $t=h+1$.